OpenStudy (anonymous):

show that \[2x ^{2}[x \times(-1)+(3-x)] -x(3-x)(4x) \over (2x ^{2})^{2}\] can reduce to \[-3 \over 2x ^{2}\]

OpenStudy (anonymous):

\[\frac{2x^2(x (-1)+(3-x))-x(3-x)4x}{\left(2x^2\right)^2} \]\[\frac{6 x^2-4 x^3-12 x^2+4 x^3}{\left(4 x^4\right)} \]\[\frac{-6 x^2}{\left(4 x^4\right)} \]\[-\frac{3}{2 x^2} \]

OpenStudy (anonymous):

how did you go from steps 1-2? like the top part?

OpenStudy (anonymous):

Work from the inside out. Multiply x*(-1) -> -x +3 -x is -2x+3 2x^2(-2x+3)=-4x^3+6x^2 Over at the right side -x(3-x)4x is (-3x+x^2)4x = -12x^2+4x^3 Putting it all together, -4x^3 + 6x^2 -12x^2 + 4x^3 = -12x^2 The -4x^3 and +4x^3 subtract out. 6x^2 - 12x^2 is equal to -6x^2

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

Your welcome.